A person has $1500.00 to invest for 1 year. He has a choice of a credit union account that pays 4.5% annual simple interest and a bank account that pays 4% interest compounded quarterly. Detemine which account would pay more interest and by how much?
Answer by stanbon(48554)

Simple interest Formula: Interest=Prt=1500*0.045(1)=$67.50
Compound interest Formula: Interest =P((1+r/n)^(nt) -1}
Interest=1500[(1+0.045/4)^4 -1]
=1500[(1.01125)^4 - 1]
=1500[0.045765...]
=$68.65
68.65-67.50=$1.15
Compound interest is better by $1.15 the first year.
Cheers,
Stan H.